# Invisible neutron decay and light BSM particles

### v1

@@ I Introduction
Searches for nucleon decays by Super-Kamiokande have pushed limits to $10^{34}$ yr and above for several decay channels. The best known examples are probably $p\to\ell^{+}\pi^{0}$ ($\ell=e,\mu$) Super-Kamiokande:2020wjk and $n\to\bar{\nu}K^{0}$ Super-Kamiokande:2025ibz , which are motivated by Grand Unified Theories (GUTs). Less known are decays to three charged leptons, where Super-Kamiokande also gives limits in excess of $10^{34}$ yr now Super-Kamiokande:2020tor . Limits on many other channels, however, are less stringent, sometimes by orders of magnitude, see Takhistov:2016eqm ; ParticleDataGroup:2024cfk . On the other hand, nucleon decay searches form an important part of the agenda of next generation underground detectors such as JUNO JUNO:2015zny , Hyper-Kamiokande Hyper-Kamiokande:2018ofw or DUNE DUNE:2020lwj and significant improvements in sensitivity on many nucleon decay modes are expected from these experiments in the foreseeable future.
@@
Traditionally, searches for proton decay were motivated by the predictions from GUTs, such as the classical $SU(5)$ Georgi:1974sy , for a review see for example Nath:2006ut . More recently, however, many theoretical papers have taken a more agnostic view and studied nucleon decays from the point of view of effective field theory (EFT) deGouvea:2014lva ; Kobach:2016ami ; Hambye:2017qix ; Fonseca:2018ehk ; Helo:2018bgb ; Helo:2019yqp ; Heeck:2019kgr ; Beneito:2023xbk ; Gargalionis:2024nij ; Heeck:2024jei ; Li:2024liy ; Li:2025slp ; IBeneito:2025nby ; Heeck:2025btc ; Heeck:2025uwh . [^1] Here we also cite some pioneering works on nucleon decays in SMEFT Weinberg:1979sa ; Wilczek:1979hc ; Weinberg:1980bf ; Weldon:1980gi ; Abbott:1980zj ; Claudson:1981gh and the studies on the decompositions of the relevant effective operators Bowes:1996xy ; Kovalenko:2002eh ; Arnold:2012sd ; Assad:2017iib ; deBlas:2017xtg ; Li:2023cwy . In Standard Model Effective Field Theory (SMEFT), one can classify different decay modes according to the dimensionality of the corresponding operators. For example, two-body decays with $\Delta(B+L)=2$ are described by $d=6$ operators, $\Delta(B-L)=2$ are $d=7$, genuine three-body decays need at least $d=9$ operators Heeck:2019kgr .
In the current paper, we are interested in invisible neutron decay. Considering only SM particles, the only possibility for a neutron to decay into “nothing” is the channel $n\to 3\nu$. [^2] Here both $\nu$ and $\bar{\nu}$ are mentioned as $\nu$. SMEFT operators with three leptons have been discussed in Hambye:2017qix . In SMEFT, the lowest dimensional operator with three leptons has $d=9$, one example is ${\cal O}\propto(QL)({\bar{L}}{\bar{L}})(d_{R}d_{R})$. Ref. Hambye:2017qix lists a total of 16 operators of this type at $d=9$. Some of these $d=9$ operators can generate the decay $n\to 3\nu$. However, all operators that can do so, always also generate decays to charged leptons with very similar rates. Since nucleon decay searches are much more sensitive to final states with charged leptons Super-Kamiokande:2020tor than to $n\to 3\nu$ ParticleDataGroup:2024cfk , $d=9$ operators can not generate invisible neutron decay at observable rates.
@@
In SMEFT one has to go up to $d=12$, to find for the first time an operator capable of generating invisible neutron decay, without the corresponding decays to charged leptons, as already realized by Weinberg more than 40 years ago Weinberg:1980bf . Weinberg gave as an example the operator Weinberg:1980bf :
@@
$\mathscr{L}^{d=12}_{\rm W}=\frac{c_{\text{W}}}{\Lambda^{8}}\epsilon_{IJK}\epsilon^{ij}\epsilon^{kl}\epsilon^{mn}\left(\overline{u_{R}}^{I}L_{i}\right)\left(\overline{d_{R}}^{J}L_{k}\right)\left(\overline{d_{R}}^{K}L_{m}\right)H_{j}H_{l}H_{n},$ (1)
@@
where $\Lambda$ is the typical scale of the origin of this effective operator and $c_{\text{W}}$ is the Wilson coefficient. In Eq. (1) we have suppressed generation indices for simplicity, $\epsilon_{IJK}$ and $\epsilon^{ij}$ are anti-symmetric tensors for the contraction of $SU(3)_{c}$ and $SU(2)_{L}$ indices. This operator leads to the decay mode $n\to 3\nu$. For this $d=12$ operator one can estimate the nucleon decay width for the 3-body final state roughly as Heeck:2019kgr
@@
Here, $\beta_{h}\simeq 0.014$ GeV3 is the nuclear matrix element Aoki:2017puj , $m_{N}$ the nucleon mass, $\langle H^{0}\rangle\simeq 174$ GeV the vacuum expectation value of the SM Higgs field, and $c_{\text{W}}$ is taken to be unity.
@@
For invisible neutron decay the particle data booklet ParticleDataGroup:2024cfk cites an old Kamiokande result (from 1993) Kamiokande:1993ivj ($4.9\cdot 10^{26}$ yr), but there are more recent and stronger bounds set by SNO SNO:2003lol ($2\cdot 10^{29}$), KamLAND KamLAND:2005pen ($5.8\cdot 10^{29}$ yr, cf. also Kamyshkov:2002wp ) and two analysis by the SNO+ collaboration SNO:2018ydj ($2.5\cdot 10^{29}$ yr) and SNO:2022trz ($9.0\cdot 10^{29}$ yr). For the nearer and farther future sensitivity estimates have been given by JUNO JUNO:2024pur ($5\cdot 10^{31}$ yr) and for THEIA (80kt), that could ultimately reach $\sim 4\cdot 10^{32}$ yr according to their proposal Theia:2019non .
@@
However, there exists a simple possibility to generate invisible neutron decay with lower dimensional operators: We can extend the SM particle content with new light degrees of freedom. In this paper we discuss four different possibilities. First we add (i) (at least two) right-handed neutrinos $N_{R}$, also sometimes called sterile neutrinos in the literature, cf. Acero:2022wqg ; Dasgupta:2021ies ; Diaz:2019fwt ; Boser:2019rta ; Dentler:2018sju ; Kopp:2013vaa ; Abazajian:2012ys for reviews and current bounds, motivated from various neutrino experiments LSND:2001aii ; MiniBooNE:2010idf ; MiniBooNE:2020pnu ; Elliott:2023cvh ; Barinov:2021asz ; Dentler:2017tkw ; STEREO:2022nzk . In addition, the introduction of such light degrees of freedom is often discussed in the context of cosmology, e.g., the stringent bound on neutrino masses from recent cosmological observations DESI:2024mwx ; DESI:2024hhd ; DESI:2025zgx ; DESI:2025ejh maybe be alleviated by the mixing between active neutrinos and massless sterile neutrinos Farzan:2015pca ; Escudero:2022gez ; Benso:2024qrg ; Ota:2024byu . In $N_{R}$SMEFT, for a list of basis operators see Bhattacharya:2015vja ; Liao:2016qyd ; Li:2021tsq , invisible neutron decay operators appear at $d=9$.
@@
Second, we add (ii) a right-handed neutrino and an axion-like particle (ALP), $a$. The axion is one of the best-motivated light new physics fields Weinberg:1977ma ; Wilczek:1977pj . It arises as a consequence of the breaking of the Peccei-Quinn symmetry introduced to solve the strong CP problem Peccei:1977hh ; Peccei:1977ur . For experimental searches and phenomenology of ALPs, cf. e.g., Graham:2015ouw ; Brivio:2017ije ; Bauer:2017ris ; Irastorza:2018dyq ; Bauer:2018uxu ; Gavela:2019wzg ; Merlo:2019anv ; Alda:2025uwo and for cosmological aspects of ALPs, cf. e.g., Cadamuro:2011fd ; Arias:2012az ; Marsh:2015xka ; Dror:2021nyr ; Chadha-Day:2021szb ; Adams:2022pbo . The basis operators of aaSMEFT, the effective theory with the SM particle content plus an ALP, are listed in Song:2023lxf ; Song:2023jqm ; Grojean:2023tsd . In $aN_{R}$SMEFT, i.e. SMEFT with $N_{R}$s and ALPs, invisible neutron decay operators appear at d=8d=8.333Basis operators with a given particle content can be constructed up to a given mass dimension $d$ by AutoEFT Harlander:2023ozs or also Sym2Int Fonseca:2017lem
@@
This introduction would not be complete without mentioning that invisible neutron decay has been discussed in some earlier papers. For example, in Dvali:1999hn the authors speculated about “decays to nothing” in models with extra time dimensions. The authors of Mohapatra:2002ug considered universal extra dimensions and an extended gauge group, to have the neutron decay dominantly into $n\to\nu\bar{\nu}_{s}\bar{\nu}_{s}$, where $\nu_{s}$ is some sterile neutrino. The effective operators relevant to $n\rightarrow\bar{\nu}_{s}\bar{\nu}^{\prime}_{s}\bar{\nu}^{\prime\prime}_{s}$ are listed in Girmohanta:2019fsx in the context of an extra dimension model. In Girmohanta:2020eav invisible neutron decay was studied in a left-right symmetric model with large extra dimensions. Some recent works have considered also the possibility of a “dark” (invisible) neutron decay Fornal:2018eol ; Barducci:2018rlx ; Cline:2018ami ; Berezhiani:2018udo ; Fornal:2020gto in relation to the neutron life-time anomaly Czarnecki:2018okw . And, finally, a particular UV model for invisible di-neutron decay to two singlet fermions has been studied in  Hao:2022xyj , where the authors discussed also possible constraints from neutron stars.
@@ II Setup: Models and effective operators
In this section we will describe the theoretical setup. We consider four different extensions of the SM particle content. For each case, we briefly describe the model. We then define the operators relevant for invisible neutron decay. We give an estimate of the neutron decay half-live in each case. Finally, in an additional subsection we discuss a possible caveat. The decomposition of the operators, defined here, is discussed in section 3.
@@ II.1 Variant I: Right-handed neutrinos
Massive neutrinos can be either Dirac or Majorana particles. In the classical type-I seesaw model Minkowski:1977sc ; Yanagida:1979as ; Mohapatra:1979ia ; Schechter:1980gr right-handed neutrinos have a Majorana mass term and the Lagrangian can be written as:
@@
Here, $\alpha,\beta,\gamma$ are SM generation indices, while $i,j,k$ runs over the generation of $N_{R}$. To keep the expressions simpler, we have suppressed $SU(3)_{c}$ (and $SU(2)_{L}$) indices. Here, and in all cases discussed below the colour of the three quarks must be contracted with a completely anti-symmetric tensor, $\epsilon_{IJK}$. It is straightforward to show that both of these operators vanish in case there is only one copy of $N_{R}$, we have checked this also with Sym2Int Fonseca:2017lem .
@@
Here again, $\beta_{h}\simeq 0.014$ GeV3 is the nuclear matrix element Aoki:2017puj and the half-live is normalized to the current bound from SNO+ SNO:2022trz . We can apply Eq. (6) to the different existing and projected half-life limits and obtain estimates for the sensitivity to the new physics scale $\Lambda$:
@@ II.2 Variant II: A right-handed neutrino and a ALP
Axion-like particles (ALPs) appear in many extensions of the SM. While the ALP mass is a free parameter, the couplings of the ALPs with SM fields are protected by an approximate classical shift symmetry, as is the case for the classical axion. Many different studies of properties of and searches for ALPs have been published, see for example Brivio:2017ije ; Bauer:2021mvw ; Bauer:2017ris ; Biekotter:2025fll . Here we only briefly mention that ALPs could also explan the dark matter Jaeckel:2010ni ; Alekhin:2015byh and/or be long-lived particles Alekhin:2015byh ; Alimena:2019zri .
The ALP Lagrangian up to $d=5$ contains the following terms Georgi:1986df

### v2

@@ I Introduction
Searches for nucleon decays by Super-Kamiokande have pushed limits to $10^{34}$ yr and above for several decay channels. The best known examples are probably $p\to\ell^{+}\pi^{0}$ ($\ell=e,\mu$) Takenaka and others (2020) and $n\to\bar{\nu}K^{0}$ Yamauchi and others (2025), which are motivated by Grand Unified Theories (GUTs). Less known are decays to three charged leptons, where Super-Kamiokande also gives limits in excess of $10^{34}$ yr now Tanaka and others (2020). Limits on many other channels, however, are less stringent, sometimes by orders of magnitude, see Takhistov (2016); Navas and others (2024). On the other hand, nucleon decay searches form an important part of the agenda of next generation underground detectors such as JUNO An and others (2016), Hyper-Kamiokande Abe and others (2018) or DUNE Abi and others (2020) and significant improvements in sensitivity on many nucleon decay modes are expected from these experiments in the foreseeable future.
@@
Traditionally, searches for proton decay were motivated by the predictions from GUTs, such as the classical $SU(5)$ Georgi and Glashow (1974), for a review see for example Nath and Fileviez Perez (2007). More recently, however, many theoretical papers have taken a more agnostic view and studied nucleon decays from the point of view of effective field theory (EFT) de Gouvea et al. (2014); Kobach (2016); Hambye and Heeck (2018); Fonseca et al. (2018); Helo et al. (2018, 2019); Heeck and Takhistov (2020); Beneito et al. (2024); Fridell et al. (2024); Gargalionis et al. (2024); Heeck and Watkins (2024); Li et al. (2024a, 2025); I Beneito et al. (2025); Heeck and Sokhashvili (2025); Heeck and Shoemaker (2025); Fan et al. (2025). [^1] Here we also cite some pioneering works on nucleon decays in SMEFT Weinberg (1979); Wilczek and Zee (1979); Weinberg (1980); Weldon and Zee (1980); Abbott and Wise (1980); Claudson et al. (1982) and the studies on the decompositions of the relevant effective operators Bowes et al. (1996); Kovalenko and Schmidt (2003); Arnold et al. (2013); Assad et al. (2018); de Blas et al. (2018); Li et al. (2024b). In Standard Model Effective Field Theory (SMEFT), one can classify different decay modes according to the dimensionality of the corresponding operators. For example, two-body decays with $\Delta(B+L)=2$ are described by $d=6$ operators, $\Delta(B-L)=2$ are $d=7$, genuine three-body decays need at least $d=9$ operators Heeck and Takhistov (2020).
@@
In the current paper, we are interested in invisible neutron decay. Considering only SM particles, the only possibility for a neutron to decay into “nothing” is the channel $n\to 3\nu$. [^2] Here both $\nu$ and $\bar{\nu}$ are mentioned as $\nu$. In SMEFT one has to go up to $d=12$, to find for the first time an operator capable of generating invisible neutron decay, without the corresponding decays to charged leptons, as already realized by Weinberg more than 40 years ago Weinberg (1980). Weinberg gave as an example the operator Weinberg (1980):
@@
$\mathscr{L}_{d=12}^{\rm W}=\frac{c_{12}^{\text{W}}}{\Lambda^{8}}\epsilon_{IJK}\epsilon^{ij}\epsilon^{kl}\epsilon^{mn}\left(\overline{u_{R}}^{I}L_{i}\right)\left(\overline{d_{R}}^{J}L_{k}\right)\left(\overline{d_{R}}^{K}L_{m}\right)H_{j}H_{l}H_{n},$ (1)
@@
where $\Lambda$ is the typical scale of the origin of this effective operator and $c_{12}^{\text{W}}$ is the Wilson coefficient. In Eq. (1) we have suppressed generation indices for simplicity, $\epsilon_{IJK}$ and $\epsilon^{ij}$ are anti-symmetric tensors for the contraction of $SU(3)_{c}$ and $SU(2)_{L}$ indices. For this $d=12$ operator one can estimate the nucleon decay width for the 3-body final state $n\to 3\nu$ roughly as Heeck and Takhistov (2020)
@@
Here, $\beta_{h}\simeq 0.014$ GeV3 is the nuclear matrix element Aoki et al. (2017), $m_{N}$ the nucleon mass, $\langle H^{0}\rangle\simeq 174$ GeV the vacuum expectation value of the SM Higgs field, and $c_{12}^{\text{W}}$ is taken to be unity.
Since the operator in Eq. (1) by itself will not generate final states with charged leptons at any measurable level, one might be tempted to assume that just having a low enough scale $\Lambda$ would be sufficient to create an observable rate for the invisible neutron decay. However, this statement neglects that there are lower dimensional operators that will generate neutron decays with charged lepton final states. SMEFT operators with three leptons have been discussed in Hambye and Heeck (2018). In SMEFT, the lowest dimensional operator with three leptons has $d=9$, one example is ${\cal O}\propto(QL)({\bar{L}}{\bar{L}})(d_{R}d_{R})$. Ref. Hambye and Heeck (2018) lists a total of 16 operators of this type at $d=9$. Some of these $d=9$ operators can generate the decay $n\to 3\nu$. However, all the operators that can do so, always also generate decays to charged leptons with very similar rates. Since nucleon decay searches are much more sensitive to final states with charged leptons Tanaka and others (2020) than to $n\to 3\nu$ Navas and others (2024), none of the $d=9$ operators can generate invisible neutron decay at observable rates. Moreover, if any of these $d=9$ operators is present in the theory, one expects it to dominate the nucleon decay rate by many orders of magnitude w.r.t the $d=12$ operator in Eq. (1).
To avoid this conclusion, one could argue that all $d=9$ operators are actually forbidden by some symmetry, such that the $d=12$ operator in Eq. (1) is the lowest dimension at which neutron decay is possible. However, this is not possible. While some specific symmetry could indeed eliminate all $d=9$ operators, at $d=10$ one finds an operator with three leptons Weinberg (1980): ${\cal O}\propto d^{c}d^{c}d^{c}LLLH$. [^3] This operator is non-zero only for two or more generations of down quarks or leptons. This operator induces the decay $n\to e^{-}\pi^{+}\nu\nu$ Fonseca et al. (2018) and will dominate the decay rate, unless its Wilson coefficient is fine-tuned to be very close to zero. As we show in appendix B one cannot forbid this operator by any symmetry, if the $d=12$ operator is to be allowed. Thus, within pure SMEFT, one expects that invisible neutron decay is always accompanied by neutron decays with charged lepton final states.
@@
For invisible neutron decay the particle data booklet Navas and others (2024) cites an old Kamiokande result (from 1993) Suzuki and others (1993) ($4.9\cdot 10^{26}$ yr), but there are more recent and stronger bounds set by SNO Ahmed and others (2004) ($2\cdot 10^{29}$), KamLAND Araki and others (2006) ($5.8\cdot 10^{29}$ yr, cf. also Kamyshkov and Kolbe (2003)) and two analysis by the SNO+ collaboration Anderson and others (2019) ($2.5\cdot 10^{29}$ yr) and Allega and others (2022) ($9.0\cdot 10^{29}$ yr). For the nearer and farther future sensitivity estimates have been given by JUNO Abusleme and others (2025) ($5\cdot 10^{31}$ yr) and for THEIA (80kt), that could ultimately reach $\sim 4\cdot 10^{32}$ yr according to their proposal Askins and others (2020).
@@
However, there exists a simple possibility to generate invisible neutron decay with lower dimensional operators and without generating sizeable rates into final states with charged leptons: We can extend the SM particle content with new light degrees of freedom. In this paper we discuss four different possibilities. First we add (i) (at least two) right-handed neutrinos $N_{R}$, also sometimes called sterile neutrinos in the literature, cf. Acero and others (2024); Dasgupta and Kopp (2021); Diaz et al. (2020); Böser et al. (2020); Dentler et al. (2018); Kopp et al. (2013); Abazajian and others (2012) for reviews and current bounds, motivated from various neutrino experiments Aguilar and others (2001); Aguilar-Arevalo and others (2010, 2021); Elliott et al. (2024); Barinov and others (2022); Dentler et al. (2017); Almazán and others (2023). In addition, the introduction of such light degrees of freedom is often discussed in the context of cosmology, e.g., the stringent bound on neutrino masses from recent cosmological observations Adame and others (2025a, b); Abdul Karim and others (2025); Elbers and others (2025) maybe be alleviated by the mixing between active neutrinos and massless sterile neutrinos Farzan and Hannestad (2016); Escudero et al. (2023); Benso et al. (2025); Ota (2025). In $N_{R}$SMEFT, for a list of basis operators see Bhattacharya and Wudka (2016); Liao and Ma (2017); Li et al. (2021), invisible neutron decay operators appear at $d=9$.
@@
Second, we add (ii) a right-handed neutrino and an axion-like particle (ALP), $a$. The axion is one of the best-motivated light new physics fields Weinberg (1978); Wilczek (1978). It arises as a consequence of the breaking of the Peccei-Quinn symmetry introduced to solve the strong CP problem Peccei and Quinn (1977b, a). For experimental searches and phenomenology of ALPs, cf. e.g., Graham et al. (2015); Brivio et al. (2017); Bauer et al. (2017); Irastorza and Redondo (2018); Bauer et al. (2019); Gavela et al. (2019); Merlo et al. (2019); Alda et al. (2025) and for cosmological aspects of ALPs, cf. e.g., Cadamuro and Redondo (2012); Arias et al. (2012); Marsh (2016); Dror et al. (2021); Chadha-Day et al. (2022); Adams and others (2022). The basis operators of aaSMEFT, the effective theory with the SM particle content plus an ALP, are listed in Song et al. (2024b, a); Grojean et al. (2023). In $aN_{R}$SMEFT, i.e. SMEFT with $N_{R}$s and ALPs, invisible neutron decay operators appear at d=8d=8.444Basis operators with a given particle content can be constructed up to a given mass dimension $d$ by AutoEFT Harlander and Schaaf (2024) or also Sym2Int Fonseca (2017)
@@
This introduction would not be complete without mentioning that invisible neutron decay has been discussed in some earlier papers. For example, in Dvali et al. (1999) the authors speculated about “decays to nothing” in models with extra time dimensions. The authors of Mohapatra and Perez-Lorenzana (2003) considered universal extra dimensions and an extended gauge group, to have the neutron decay dominantly into $n\to\nu\bar{\nu}_{s}\bar{\nu}_{s}$, where $\nu_{s}$ is some sterile neutrino. The effective operators relevant to $n\rightarrow\bar{\nu}_{s}\bar{\nu}^{\prime}_{s}\bar{\nu}^{\prime\prime}_{s}$ are listed in Girmohanta and Shrock (2020) in the context of an extra dimension model. In Girmohanta (2021) invisible neutron decay was studied in a left-right symmetric model with large extra dimensions. Some recent works have considered also the possibility of a “dark” (invisible) neutron decay Fornal and Grinstein (2018); Barducci et al. (2018); Cline and Cornell (2018); Berezhiani (2019); Fornal and Grinstein (2020) in relation to the neutron life-time anomaly Czarnecki et al. (2018). And, finally, a particular UV model for invisible di-neutron decay to two singlet fermions has been studied in  Hao and Ni (2023), where the authors discussed also possible constraints from neutron stars.
@@ II Setup: Models and effective operators
In this section we will describe the theoretical setup. We consider four different extensions of the SM particle content. For each case, we briefly describe the model. We then define the operators relevant for invisible neutron decay. We give an estimate of the neutron decay half-life in each case. Finally, in an additional subsection we discuss a possible caveat. The decomposition of the operators, defined here, is discussed in section 3.
@@ II.1 Variant I: Right-handed neutrinos
Massive neutrinos can be either Dirac or Majorana particles. In the classical type-I seesaw model Minkowski (1977); Yanagida (1979); Mohapatra and Senjanovic (1980); Schechter and Valle (1980) right-handed neutrinos have a Majorana mass term and the Lagrangian can be written as:
@@
Here, $\alpha,\beta,\gamma$ are SM generation indices, while $i,j,k$ runs over the generation of $N_{R}$. To keep the expressions simpler, we have suppressed $SU(3)_{c}$ (and $SU(2)_{L}$) indices. Here, and in all cases discussed below the colour of the three quarks must be contracted with a completely anti-symmetric tensor, $\epsilon_{IJK}$. It is straightforward to show that both of these operators vanish in case there is only one copy of $N_{R}$, we have checked this also with Sym2Int Fonseca (2017).
@@
Here again, $\beta_{h}\simeq 0.014$ GeV3 is the nuclear matrix element Aoki et al. (2017) and the half-life is normalized to the current bound from SNO+ Allega and others (2022). We can apply Eq. (6) to the different existing and projected half-life limits and obtain estimates for the sensitivity to the new physics scale $\Lambda$:
@@ II.2 Variant II: A right-handed neutrino and a ALP
Axion-like particles (ALPs) appear in many extensions of the SM. While the ALP mass is a free parameter, the couplings of the ALPs with SM fields are protected by an approximate classical shift symmetry, as is the case for the classical axion. Many different studies of properties of and searches for ALPs have been published, see for example Brivio et al. (2017); Bauer et al. (2022, 2017); Biekötter and Mimasu (2025). Here we only briefly mention that ALPs could also explan the dark matter Jaeckel and Ringwald (2010); Alekhin and others (2016) and/or be long-lived particles Alekhin and others (2016); Alimena and others (2020).
The ALP Lagrangian up to $d=5$ contains the following terms Georgi et al. (1986)